OpenStudy (anonymous):

Events A and B are overlapping. P(A) =6/17 ; P(B) =5/17 ; P(A and B) =3/17 Find P(A or B). 14/17 8/17 11/17 9/17

OpenStudy (neer2890):

p(A or B)= p(A)+p(B)-p(A and B) so p(A or B)=6/17+5/17-3/17=8/17

OpenStudy (dan815):

p(a)+p(b)-p(a+b)=p(a or b)